# What could go wrong? Discovering and describing failure modes in computer vision

### v1

@@ Appendix A Datasets and Tasks
Urban scene segmentation. We consider a ConvNeXt [29] segmentation model trained on Cityscapes [3], comprised of images from 50 European cities collected at daytime in clear weather conditions. We build our evaluation sets ${\mathcal{X}}$ using three datasets: i) WildDash 2 (WD2) [52], which contains challenging visual conditions such as motion blur, various road types, difficult weather, etc.  and ii) India Driving Dataset (IDD) [45], which contains images from Hyderabad, Bangalore and their peripheries, iii) ACDC [37] contains images recorded in adverse conditions, namely fog, rain, snow and night representing challenging domain shift for the Cityscapes model. For the main experiments we use the user-defined sentence set $\mathcal{S}$ with 148 sentences describing content related to urban scenes, road conditions, and image quality (described in Appendix C). In Appendix D, we show further results on WD2 with both a smaller GT-based sentence set derived from the metadata information available with the dataset and a larger sentence set of 1016 sentences generated automatically by an LLM (both detailed in Appendix C).
@@
We consider two cases, an unsupervised and a supervised one. In the former case we consider the full set and split it with entropy values computed on class predictions. In the latter case, similarly to [9, 51], we consider each class separately and use the class probability scores to split the data into easy and hard sets.
@@
ImageNet-1K classification. We consider three architectures (ResNet-50 [12], VIT-B-16 [7], and EfficientNet-B1 [43]) trained on ImageNet-1K [36] and evaluate these models on the validation set. We build ${\mathcal{S}}$ as the union of a set of sentences {“An image of a <class>”} (for the 1K classes) with a set of sentences corresponding to place classes from Places365 [56] {“An image taken at <location>”}, and a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc.  In total we have 1417 sentences (see details in Appendix C). To assess the relevance between a sentence and an image, we use a combination of ground truth based information (the ones related to the 1000 class names) and a VQA model to assess the relevance for the other sentences (see Sec. C.3).
@@
In particular, we focus on the top-1 classification performance of ImageNet models and aim at analyzing failure cases for this task—namely, the reasons for not predicting the correct class on top. In ImageNet classification, several errors arise from confusions between similar classes. For example, the model will miss-classify dog breeds, bird/fish species, similar objects, etc. In such cases, knowing the correct class label can help to deduce the main reason from the retrieved description. Indeed, for example, if the selected sentence is {“An image of a water snake”} without the knowledge of the correct class, it is difficult to interpret whether the model predicts false positives (images of different type of snakes) or the model failed to recognize the water snake (false negative) for some reasons. Furthermore, while, class independent failure reasons such as those related to image capturing conditions (e.g.blurred or low contrast images) might emerge when processing the full dataset as a whole, errors related to the image content is more difficult to interpret if it is a failure reason without the target class (e.g.finding that the image was taken at a beach has different effect on the model for the swimming suit class than for the cow or ski suit class). Finally, even if we use a very large number of clusters with the full set, the content of the nearest easy cluster might differ in many aspects from the hard one, making it difficult to select the actual failure reason.
@@
For all these reasons we opted to follow the prior art and apply LBEE methods by analyzing the data for each class independently. Still, in contrast to prior art, where only false negative errors are considered (the images that belong to the given class but labeled with a wrong class label), we also include false positives. The easy set ${\mathcal{X}}^{e}$ contains the images where the class has been correctly predicted (images from the current class).
Appendix B Comparison with prior art
[…]
In our case, there is no learning involved: we simply split and cluster the data as discussed in Sec. 4. Then, for a given ($A,B$) pair of sets, where $A,B\in\{{\textit{Biased}},{\textit{Unbiased}}\}$, we assign images from the set $B$ to the cluster prototypes obtained with the images in set $A$, where the union of easy and hard prototypes is considered. We use the class probability to split the data—the same measure used by FACT to learn the partitioning function.
Evaluation protocol. Before discussing the comparative results, we recall the metrics and the evaluation protocol we use, proposed by [9, 51]. In particular they propose to rely on the Precision-at-K (P and K ) to measure how accurately a partitioning method aggregates samples from the same context given a class. Formally, let ${\mathcal{X}}$ denote the test set, $\{{\bf z}_{n}\}_{n=1}^{N}$ the ground-truth partitions and $\{\widehat{{\bf z}}_{m}\}_{m=1}^{M}$ the predicted set of discovered partitions. Given $K$, for each ground-truth partition ${\bf z}_{n}$ the most "similar" predicted partition $\widehat{{\bf z}}_{m}$ is selected based on the size of the intersection ${\bf z}_{n}\cap\widehat{{\bf z}}^{K}_{m}$, where $\widehat{{\bf z}}^{K}_{m}$ is the subset containing the top-$K$ elements of ${\mathcal{X}}$ according to the likelihood of belonging to $\widehat{{\bf z}}_{m}$. For FACTS/DOMINO the ranking corresponds to the partitioning function’s output, while in our case to the distance to the prototype. If we denote by $\widehat{{\bf z}}^{K}_{nm}$ the set that has the largest overlap with ${\bf z}_{n}$ given $K$, the P and K is defined as
[…]
where the average is computed, as in [51], over the five bias-conflicting groups (excluding the spuriously correlated GT partitions, such as waterways in water).
@@ Appendix B Comparison with prior art
Experimental results. We show in Tab. 3 the P and K values averaged over the six classes for the different NICO++{++}+ + sets. Each time we show results averaged over three different variants of the current model [^8] Since we were not able to reproduce the original paper’s numbers with the code, namely, the results corresponding to the (Biased,Unbiased) pairs, for fairness we ran their code with three different seeds and provide here the averaged results. We do not show the variances in the table to keep it simple, but the values are most often between 1 and 2.. In our case, we used the exact same ResNet50 models and the same Biased and Unbiased sets on which we perform the split and the clustering. We use five easy and five hard clusters as it is our default setting for the per-class case and as in the code from [51] the default number of slices was also set to ten.
From Tab. 3 we can see that our split-and-cluster method provides competitive, or sometimes even better, slice discovery performance than the more complex partition learning methods. Our method performs slightly worse than FACTS on the Unbiased set both when their partitioning was trained on the Unbiased or on the Biased set. Surprisingly though, both FACTS and DOMINO underperform when tested on the Biased set even when the partitioning function was trained on this set. In contrast, our method is more robust and performs similarly on both sets. When the FACTS model is trained on the Biased set, which strongly relies on priors about class-context associations (the configuration shown in their paper), it outperforms our model, for which the P and K decreases as the amount of class-context correlation increases (bottom part of Tab. 3). We hypothesize that the reason behind this result is that we do not have enough data for a proper clustering (see Tab. 2), while FACTS manages to learn the partitioning function even with a a few samples by exploiting the strong correlation between the classes and dominant context. As the model is trained mainly with images with specific combinations of class and correlated context—for example, mammals in rocky environments—the model rarely sees images from other classes in this context. Consequently, when the model learns to predict a class $c$" with high probability, it also associates this high probability with the presence of the corresponding context (e.g.rocks) even if the class (mammals) is absent. Therefore, when analyzing another class, such as plants, and learn the partitioning based on all six probabilities, the partitioning function primarily learns to assigns images based on the highest probability score, which is an indicator of the context. For instance, a plant image taken in a rocky environment will have a high probability score for the mammal class, indicating the presence of rocks. Note that in our case, when analyzing the plant class, we disregard the probabilities of other classes, thereby using much less prior information about the image context.
@@ Appendix C Defining the sentence sets
Our paper focuses on two objectives, 1) identifying error modes in computer vision models and 2) describing them with natural language. To achieve this, we rely on a large set of sentences (${\mathcal{S}}$), that can either be provided or automatically generated. This sentence set is expected to cover various potential reasons for model failures given a task. In this section, we first discuss three approaches to construct such sentence sets (Sec. C.1); then we detail how we create the ground-truth sentence set ${\mathcal{S}}^{*}_{\beta}$ for LBEE (Sec. C.2); finally, we show how to obtain pseudo-GT relevance scores between images and sentences using Visual-Question Answering (VQA) tools (Sec. C.3).
@@ C.1 Building the sentence set ${\mathcal{S}}$
User-defined. The set of sentences can be designed ad hoc for the task at hand by the end user who has the expert knowledge and can envisage the main difficulties a model might face. While such sentence set might be non-exhaustive, it can cover the most important failure cases. In this case, our methods are able to select the ones that confirm the user’s concerns (with additional visual support).
We follow this setting to build ${\mathcal{S}}$ in our experiments on the urban scene segmentation task. More specifically, we manually defined 148 sentences describing content related to urban scenes (buildings, pedestrians, traffic, two-wheels, trees, animals, garbage, etc. ) in the form of {“An image of a <urban scene content>”}, road conditions (weather, lighting, season) in form of {“An image taken in/at <condition>”} or image quality (underexposure, motion blur, etc. ) in form of {“An image taken in/at <effect>”}. A large subset of this sentence set can be seen in Tab. 5.
In this case, to get pseudo-ground-truth relevance between images and a sentence, we resort to VQA (as described in Sec. C.3).
@@
In the case of ImageNet-1K [36], the first 1K sentences (out of 1417) were also generated from the GT information, namely the class names. We consider sentences with the form {"An image of a <class>”} to which we added sentences with the form {“An image taken in <location>”} filling <location> with the place class names from Places365 [56] and a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc.
Concerning the urban scene segmentation task, we also built a GT-based sentence set for the WD2 dataset [52]. It contains annotations for 81 classes, from which we selected 74 classes (removing the ambiguous ones most difficult for VLMs to handle, namely ‘unlabeled’, ‘ego-vehicle’, ‘overlay’, ‘out-of-roi’, ‘static’, ‘dynamic’, ‘curb terrain’). The sentences have the form {“An image showing a <class>”}, where <class> was replaced each time by one of the 74 classes selected. While this allows us to generate GT relevance scores for each image in WD2, this set limits our methods in selecting only from failure causes related to the presence of these classes in the image and hence they do not cover all the possible failure reasons that may explain the model’s behavior.
LLM-generated. We also consider a set of sentences generated by a large language model (LLM). We provided GPT-3.5 with the context and the sentence structure and prompted it to provide lists of sentences related to weather conditions, vehicles, urban objects, time of day and road conditions. For each of them we used either the sentence form {“An image taken in/at <weather/road condition or time of day>”} or {“An image showing a/an <vehicle/urban object>”}, where the template is filled by the LLM. This yields to a set of 1016 sentences, including most of the ones we design manually in the User-defined set. On the other end, the limitation of this approach is the noise in the set, which can also include sentences hardly helpful for the task at end, e.g.An image taken on an unpredictable/eclectic/dystopian road” or “An image taken in a corn maze/trick-or-treat/pajama weather”.
@@ C.2 Defining the ground-truth ${\mathcal{S}}_{\beta}^{*}$ for LBEE
To evaluate the predicted sentence set ${\mathcal{R}}_{\mathcal{S}}\subset{\mathcal{S}}$ output by an LBEE approach, we must define a set of ground-truth sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ to compare our predictions to—namely, ground-truth error mode descriptions. To derive such set, we consider the function $\omega_{\theta}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}$ that measures the model’s performance on each image, denoted by $\omega_{\theta}({\bf x}_{i})$. This function can be the accuracy in the case of a classification task, the mIoU (mean Intersection over Union) score for semantic segmentation, or any other task-specific metric. Given the set of images ${\mathcal{X}}$, the average value of this measure over the entire dataset
@@
[…]

### v2

@@ Appendix A Datasets and Tasks
Urban scene segmentation. We consider a ConvNeXt [29] segmentation model trained on Cityscapes [3], comprised of images from 50 European cities collected at daytime in clear weather conditions. We build our evaluation sets ${\mathcal{X}}$ using three datasets: i) WildDash 2 (WD2) [52], which contains challenging visual conditions such as motion blur, various road types, difficult weather, etc.  and ii) India Driving Dataset (IDD) [45], which contains images from Hyderabad, Bangalore and their peripheries, iii) ACDC [37] contains images recorded in adverse conditions, namely fog, rain, snow and night representing challenging domain shift for the Cityscapes model. For the main experiments we use the user-defined sentence set $\mathcal{S}$ with 148 sentences describing content related to urban scenes, road conditions, and image quality (described in Appendix C). In Sec. 6.1.3 we compare these results with the results obtained on WD2 when using both a smaller GT-based sentence set derived from the metadata information available with the dataset and a larger sentence set of 1016 sentences generated automatically by an LLM (both sets detailed in Appendix C).
@@
We consider two cases, an unsupervised and a supervised one. In the former case, we consider the full set and split it with entropy values computed on class predictions. In the latter case, similarly to [9, 51], we consider images from each class separately and use the class probability scores to split the data into easy and hard sets.
@@
ImageNet-1K classification. We consider three different architectures (ResNet-50 [12], VIT-B-16 [7], and EfficientNet-B1 [43]) trained on the ImageNet-1K [36] training set and evaluate these models on the ImageNet-1K validation set. We build ${\mathcal{S}}$ as the union of a set of sentences {“An image of a <class>”} (for the 1K classes) with a set of sentences corresponding to some place classes from Places365 [56] in form of {“An image taken at <location>”}, and a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc.  In total we have 1417 sentences (see details in Appendix C). To assess the relevance between a sentence and an image, we use a combination of ground truth based information (for the first 1000 sentences corresponding to the 1000 class names) and a VQA model to assess the relevance for the other sentences (as described in Sec. C.3).
@@
With this dataset, we focus on the top-1 classification performance of different models and aim at analyzing failure cases for this task—namely, the reasons for not predicting the correct class on top. In ImageNet classification, several errors arise from confusions between semantically similar classes. For example, the model will miss-classify dog breeds, bird/fish species, similar objects, etc. In such cases, knowing the correct class label can help to deduce the main reason from the retrieved description. Indeed, for example, if the selected sentence is {“An image of a water snake”} without the knowledge of the correct class, it is difficult to interpret whether the model predicts false positives (images of different type of snakes) or the model failed to recognize the water snake (false negative) for some reasons. Furthermore, while, class independent failure reasons such as those related to image capturing conditions (e.g.blurred or low contrast images) might emerge when processing the full dataset as a whole, errors related to the image content is more difficult to interpret if it is a failure reason without the knowledge of the target class (e.g.finding that the image was taken at a beach has different effect on the model for the swimming suit class than for the cow or ski suit class). Finally, even if we use a very large number of clusters with the full set, the content of the nearest easy cluster might differ in many aspects from the hard one, making rather it difficult to select in priority which is the one causing the model to fail.
@@
For all these reasons we opted to follow the prior art and apply LBEE methods by analyzing the data for each class independently. Still, in contrast to prior art, where only false negative errors are considered (the images that belong to the given class but labeled with a wrong class label), we also include false positives. Hence, the easy set ${\mathcal{X}}^{e}$ contains the images where the class has been correctly predicted (images from the current class) while the hard set is a union of false positives and false negatives.
@@
Appendix B Comparison with prior art
[…]
In our case, there is no learning involved: we simply split and cluster the data as discussed in Sec. 4. Then, for a given ($A,B$) pair of sets, where $A,B\in\{{\textit{Biased}},{\textit{Unbiased}}\}$, we assign images from the set $B$ to the cluster prototypes obtained with the images in set $A$, where the union of easy and hard prototypes is considered. We use the class probability to split the data—the same measure used by FACTS [51] to learn the partitioning function, except that we only use the probability of the target class while they use the probabilities of all classes.
Evaluation protocol. Before discussing the comparative results, we recall the metrics and the evaluation protocol we use, proposed by [9, 51]. In particular they propose to rely on the Precision-at-K (P and K ) to measure how accurately a partitioning method aggregates samples from the same context given a class. Formally, let ${\mathcal{X}}$ denote the test set, $\{{\bf z}_{n}\}_{n=1}^{N}$ the ground-truth partitions and $\{\widehat{{\bf z}}_{m}\}_{m=1}^{M}$ the predicted set of discovered partitions. Given $K$, for each ground-truth partition ${\bf z}_{n}$ the most "similar" predicted partition $\widehat{{\bf z}}_{m}$ is selected based on the size of the intersection ${\bf z}_{n}\cap\widehat{{\bf z}}^{K}_{m}$, where $\widehat{{\bf z}}^{K}_{m}$ is the subset containing the top-$K$ elements of ${\mathcal{X}}$ according to the likelihood of belonging to $\widehat{{\bf z}}_{m}$. For FACTS/DOMINO the ranking corresponds to the partitioning function’s output, while in our case to the assignment to the closest prototype. If we denote by $\widehat{{\bf z}}^{K}_{nm}$ the set that has the largest overlap with ${\bf z}_{n}$ given $K$, the P and K is defined as
[…]
where the average is computed, as in [51], over the five bias-conflicting groups (excluding the spuriously correlated GT partitions, such as the context water for waterways).
@@ Appendix B Comparison with prior art
Experimental results. We show in Tab. 3 the P and K values averaged over the six classes for the different NICO++{++}+ + sets. Each time we show results averaged over three different variants of the current model trained with different seed [^7] Since we were not able to reproduce the original paper’s numbers with the code, namely, the results corresponding to the (Biased,Unbiased) pairs, for fairness we ran their code with three different seeds and provide here the averaged results. We do not show the variances in the table to keep it simple, but the values are most often between 1 and 2.. In our case, we used the exact same ResNet50 models and the same Biased and Unbiased sets on which we perform the split and the clustering. We use five easy and five hard clusters as it is our default setting for the per-class case and as in the code from [51] the default number of slices was also set to ten.
From Tab. 3 we can see that our split-and-cluster method provides competitive, or sometimes even better, slice discovery performance than the more complex partition learning methods. Our method performs slightly worse than FACTS on the Unbiased set both when their partitioning was trained on the Unbiased or on the Biased set. Surprisingly though, both FACTS and DOMINO underperform when tested on the Biased set even when the partitioning function was trained on this set. In contrast, our method is more robust and performs similarly on both sets. When the FACTS model is trained on the Biased set, which strongly relies on priors about class-context associations (the configuration shown in their paper), it outperforms our model, for which the P and K decreases as the amount of class-context correlation increases (bottom part of Tab. 3). We hypothesize that the reason behind this result is that we do not have enough data for a proper clustering (see Tab. 2), while FACTS manages to learn the partitioning function even with a a few samples by exploiting the strong correlation between the classes and dominant context. As the model is trained mainly with images with specific combinations of class and correlated context—for example, mammals in rocky environments—the model rarely sees images from other classes in this context. Consequently, when the model learns to predict a class "$c$" with high probability, it also associates this high probability with the presence of the corresponding context (e.g.rocks) even if the class (mammals) is absent. Therefore, when analyzing another class, such as plants, and learn the partitioning based on all six probabilities, the partitioning function primarily learns to assigns images based on the highest probability score, which is an indicator of the context. For instance, a plant image taken in a rocky environment will have a high probability score for the mammal class, indicating the presence of rocks. Note that in our case, when analyzing the plant class, we disregard the probabilities of other classes, thereby using much less prior information about the image context.
@@ Appendix C Defining the sentence sets
Our paper focuses on two objectives, 1) identifying error modes in computer vision models and 2) describing them with natural language. To achieve this, we rely on a large set of sentences (${\mathcal{S}}$), that can either be provided or automatically generated. This sentence set is expected to cover various potential reasons for model failures given a task. In this section, we first discuss three approaches to construct such sentence sets (Sec. C.1); then we detail how we create the ground-truth sentence set ${\mathcal{S}}^{*}_{\beta}$ for LBEE (Sec. C.2); finally, we show how to obtain pseudo-GT relevance scores between images and sentences using Visual-Question Answering (VQA) tools when not available otherwise (Sec. C.3).
@@ C.1 Building the sentence set ${\mathcal{S}}$
User-defined. The set of sentences can be designed ad hoc for the task at hand by the end user who has the expert knowledge and can envisage the main difficulties a model might face. While such sentence set might be non-exhaustive, it can cover the most important failure causes. Our method aim to select from them the ones that confirm the user’s concerns (with additional visual support). In this case, to get pseudo-ground-truth relevance between images and a sentence, we resort to VQA (as described in Sec. C.3).
We follow this setting to build ${\mathcal{S}}$ in our experiments on the urban scene segmentation task. More specifically, we manually defined 148 sentences describing content related to urban scenes (buildings, pedestrians, traffic, two-wheels, trees, animals, garbage, etc. ) in the form of {“An image of a <urban scene content>”}, road conditions (weather, lighting, season) in form of {“An image taken in/at <condition>”} or image quality (underexposure, motion blur, etc. ) in form of {“An image with <effect>”}. A large subset of this sentence set can be seen in Tab. 5.
@@
We also built a GT-based sentence set for the WD2 dataset [52] tested in the context of urban scene segmentation task. It contains annotations for 81 classes, from which we selected 74 classes (removing the ambiguous ones most difficult for VLMs to handle, namely ‘unlabeled’, ‘ego-vehicle’, ‘overlay’, ‘out-of-roi’, ‘static’, ‘dynamic’, ‘curb terrain’). The sentences have the form {“An image showing a <class>”}, where <class> was replaced each time by one of the 74 classes selected. While this allows us to generate GT relevance scores for each image in WD2, this set limits our methods in selecting only from failure causes related to the presence of these classes in the image and hence they do not cover all the possible failure reasons that may explain the model’s behavior.
In the case of ImageNet-1K [36], the first 1K sentences (out of 1417) were also generated from the GT information, namely the class names. We consider sentences with the form “{An image of a <class>”}. This was complemented by a set of user-defined sentence set related to various places in the form {“An image taken in <location>”} where we filled the <location> with the place class names from Places365 [56]. Finally we also added a set of sentences describing image type (photo, drawing), image quality (blurry, noise, JPEG compression), weather conditions, etc. yielding to a total of 1417 sentence set used for ImageNet-1K.
LLM-generated. In the case of urban scene segmentation task, we also consider a set of sentences generated by a large language model (LLM). We provided GPT-3.5 with the context and the sentence structure and prompted it to provide lists of sentences related to weather conditions, vehicles, urban objects, time of day and road conditions. For each of them we used either the sentence form {“An image taken in/at <weather/road condition or time of day>”} or {“An image showing a/an <vehicle/urban object>”}, where the template is filled by the LLM. This yields to a set of 1016 sentences, including most of the ones we design manually in the User-defined set. On the other end, the limitation of this approach is the noise in the set, which can also include sentences hardly helpful for the task at end, e.g.An image taken on an unpredictable/eclectic/dystopian road” or “An image taken in a corn maze/trick-or-treat/pajama weather”.
@@ C.2 Defining the ground-truth ${\mathcal{S}}_{\beta}^{*}$ for LBEE
To evaluate the predicted sentence set ${\mathcal{R}}_{\mathcal{S}}\subset{\mathcal{S}}$ output by an LBEE approach, we must define a set of ground-truth sentences ${\mathcal{S}}^{*}\subset{\mathcal{S}}$ to compare our predictions to—namely, ground-truth error mode descriptions. To derive such GT set, we consider the function $\omega_{\theta}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}$ that measures the model’s performance on each image ${\bf x}_{i}$. This function can be the accuracy in the case of a classification task, the mIoU (mean Intersection over Union) score for semantic segmentation, or any other task-specific metric. Given the set of images ${\mathcal{X}}$, the average value of this measure over the entire dataset
@@
Given this value and a threshold $\beta$, we define the ground-truth sentence set ${\mathcal{S}}^{*}$, associated with failure cases of ${\mathcal{M}}_{\theta}$ on ${\mathcal{X}}$, as the set of sentences for which the mean performance across images associated with the sentence falls beneath the overall mean performance by a given margin $\beta$. To retrieve the images associated with a sentence, we define a binary function $\Gamma({\bf x}_{m},s_{n})$ that outputs 1 if the sentence $s_{n}$ describes the image ${\bf x}_{m}$ (i.e., is relevant to it), and 0 otherwise. For example, if ${\bf x}_{m}$ is an image taken at night, $\Gamma({\bf x}_{m},``\textit{Image taken at night}^{\prime\prime})=1$ whereas $\Gamma({\bf x}_{m},``\textit{Image taken at sunset}^{\prime\prime})=0$. If this relevance cannot be derived directly from ground truth information (most of the cases), we resort to VQA [1] to obtain such information for arbitrary sentences on arbitrary datasets—as detailed in Sec. C.3.